Circular No. 7356

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/2000 B3 (LINEAR) An object with unusual motion that was reported as asteroidal by

LINEAR (discovery observation below) has been found to be cometary in appearance, following posting on the NEO Confirmation Page. CCD observations by P. Kušnirák (Ondřejov, 0.65-m f/3.6 reflector) on Feb. 1 show a coma diameter of 6′′ and a faint tail in p.a. 120^{o}, and F. Zoltowski (Edgewood, NM, 0.3-m f/3.3 reflector) reports a small faint tail ∼ 30′′ long in p.a. 100^{o} and a dense coma ∼ 10′′ across.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 27.24052  6^{h}29^{m}47.^{s}84  +43^{o}42′44′′.9  18.9

The orbital elements given below were computed from the complete set of astrometry given on MPEC 2000-C06:

T = 2000 Feb. 14.533 TT  ω = 130.^{o}760
e = 0.57590  Ω = 352.128 }2000.0
q = 1.69866 AU  i = 11.120

a = 4.00531 AU n^{o} = 0.122956 P = 8.02 years

SUPERNOVA 1999ct AND NO SUPERNOVA 1999cv A. Gal-Yam and D. Maoz, Tel Aviv University; and P. Guhathakurta,

University of California at Santa Cruz, report: “A spectrum of SN 1999ct (cf. IAUC 7210), obtained with the Keck-II telescope (+ LRIS spectrograph, range 380–880 nm) on 1999 Aug. 6–7 by Guhathakurta reveals that SN 1999ct is of type Ia, then ∼ 2 months past maximum brightness. The derived redshift of 0.18 is consistent with that of Abell 1697 (0.183). A spectrum of SN 1999cv (ibid.), obtained by Guhathakurta at the same time with the same instrument, is that of a typical active galactic nucleus at redshift 0.797. Broad emission lines of C IV (190.9 nm), Mg II (279.8 nm), and Hβ (486.1 nm), and narrow lines of O III (495.9 and 500.7 nm), are observed. We therefore conclude that this object is not a supernova.”

SUPERNOVA 2000C IN NGC 2415 Unfiltered CCD magnitudes: Jan. 21.198 UT, 15.00 (M. Suárez, Las

Palmas, Canary Islands); 25.697, 14.8 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan).

2000 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green